# Quantum Rainbow Codes

### v2

@@ Appendix B Unfolding of rainbow codes
An important characteristic of $D$-dimensional colour codes is the possibility to map them into $D$ copies of the toric code through a local Clifford unitary [35]. This mapping, often called "unfolding" due its interpretation in two dimensions, has many useful consequences, such as the possibility to construct colour code decoders based on toric code ones [36] or the understanding of topological excitations and defects in the colour codes [37]
We prove here the existence a local Clifford unitary map between any generic rainbow code constructed from a length-2 chain complex, and the tensor product of two codes whose qubits are supported on the 1-maximal subgraphs of the corresponding flag graph. We call this mapping "unfolding" by analogy with the colour code case, and prove its existence by generalizing the techniques developed in Ref. [36]. When applying the same toolbox to rainbow codes in higher dimensions or beyond the generic case (e.g. mixed rainbow codes), those techniques seem to fail, which suggests that no such mapping exist in general. We leave this problem open for future work.
@@ B.1 General unfolding process
To prove the existence of an unfolding map, we follow the procedure outlined in Ref. [35], that we summarize now. The goal is to construct a local Clifford unitary between a stabilizer code $\mathcal{C}_{A}$ supported on qubits $\mathcal{Q}_{A}$ to a stabilizer code $\mathcal{C}_{B}$ supported on qubits $\mathcal{Q}_{B}$. For this, the idea is to partition the two qubit sets into small parts, $\{Q_{A}^{(i)}\}$ and $\{Q_{B}^{(i)}\}$, and construct a unitary $U_{i}$ for each element of the partition, such that the overall unitary operator can be written as a tensor product of all the $U_{i}$, making it local by construction. To see the effect of such unitary on the stabilizer group of $C_{A}$, we need to consider the overlap of all the stabilizers of $C_{A}$ with each qubit subset $Q_{A}^{(i)}$. If we know how all those stabilizer overlaps are mapped by all the $U_{i}$, we can determine the effect of the overall unitary on the code. The overlap of all the stabilizers of a code $\mathcal{C}$ with any given subset of qubits $Q$ forms a group, that we call the overlap group and write
@@
The following theorem (shown in [35]) characterizes an unfolding process:
@@
    $[g,g^{\prime}]=[h,h^{\prime}]$ for each $g,g^{\prime}\in\mathcal{G}(O_{A}^{(i)})$, $h=\mathcal{M}_{i}(g)$, $h^{\prime}=\mathcal{M}_{i}(g^{\prime})$.
@@
Note that the first and second conditions prove that the there exists an isomorphism between the two overlap groups, while the last condition ensures that there is unitary implementing the explicit group isomorphism given by $\mathcal{M}_{i}$. The proof of Theorem 1 heavily relies on the framework constructed in Ref. [38].
@@ B.2 Outline of the proof
    Proving that this mapping is valid by proving that the three conditions of Theorem 1 are satisfied.
@@ B.3 Number of ancilla qubits
For the base case, let’s consider a flag graph of clique cardinality $(m_{2},0,0,\ldots)$. The only type of flag graph with no cliques of size higher than $2$ are $2$-rainbow subgraphs. Since $2$-rainbow subgraphs are cycle graphs, the number of vertices is equal to the number of edges, or equivalently to the number of $2$-cliques $m_{2}$. Therefore $v=m_{2}$ and Eq. 41 is satisfied.
Let’s assume that Eq. 41 is true for any flag graph of clique cardinality $(m_{2i})_{i\geq 1}$. Let $k\geq 2$. Let’s show that Eq. 41 is also true for a flag graph $\mathcal{F}$ of clique cardinality $(m_{2},\ldots,m_{2k-2},m_{2k}+1,m_{2k+2},\ldots)$, that is, with one more clique of size $2k$. $\mathcal{F}$ has at least one $2k$-clique of size $2k$. Let’s select one, partition its vertices into $k$ pairs, and delete every edge that does not correspond to a pair of the partition. This new graph, that we call $\widetilde{\mathcal{F}}$, is a valid flag graph, with the same number of vertices as $\mathcal{F}$, but one less $2k$-clique and $k$ more $2$-cliques (corresponding to the edges that have not been deleted). Therefore, $\widetilde{\mathcal{F}}$ has clique cardinality $(m_{2}+k,\ldots,m_{2k-2},m_{2k},m_{2k+2},\ldots)$. By the induction hypothesis, we know that the number of vertices of $\widetilde{\mathcal{F}}$ is given by Eq. 41:
@@
Since the number of vertices is the same for $\widetilde{\mathcal{F}}$ and $\mathcal{F}$, we can deduce that Eq. 43 gives the correct identity and Eq. 41 is verified for $\mathcal{F}$. ∎
@@
Figure 17: Illustration of the proof of Lemma 13.(a) Flag graph with one $6$-clique, one $4$-clique, and a rainbow rank $r=4$. (b) Corresponding clique graph, with one node per clique and an edge per connection between two cliques. Rainbow subgraphs of the flag graph correspond to cycles of the clique graph.
@@
To prove this identity, let’s define the clique graph $\mathcal{C}(\mathcal{G})$ associated to $\mathcal{G}$ as a graph where each clique of $\mathcal{G}$ correspond to a vertex in $\mathcal{C}(\mathcal{G})$, and each edge of $\mathcal{G}$ connecting the vertices of two cliques becomes an edge connecting those two cliques in $\mathcal{C}(\mathcal{G})$. An example of clique graph is shown in Fig. 17. Note that the clique graph is in reality a multigraph, as multiple edges can connect a given pair of nodes. The clique graph has the following useful properties:
@@
Indeed, each $2k$-clique corresponds to a vertex of degree $2k$, and we divide the sum by two to avoid over-counting edges. Inserting this in Eq. 48, we get
@@ B.4 Mapping construction
The mapping $\mathcal{M}$ is constructed by describing how generators of $O_{A}$ are mapped into operators supported on the cliques of $\mathcal{G}$. For a $Z$ generator supported on the vertices of a clique $\mathcal{C}$, we would like to map it into a single $Z$ operator living on $\mathcal{C}$. However, we note that the product of all the $Z$ generators of $O_{A}$ gives the identity, while the corresponding single-$Z$ operators does not cancel. Therefore, such mapping would map the identity operator into a non-identity operator, and thus cannot be an isomorphism. To solve this issue, we apply the mapping described above to all but one clique. For this clique, we transform the input generator by multiplying it to an all $X$ operator (which is an element of the center of $O_{A}$), and map this new generator to the single $Z$ operator on the clique. You can find an illustration of this in Fig. 18a.
@@
For $X$ operators, the idea is to map operators supported on the vertices of an edge into an operator acting on the two cliques incident to that edge. However, we have the converse problem as for $Z$ operators: every rainbow subgraph supports two non-trivial $X$ generators that would map to the identity, namely the product of all the $c$-colored edges for each choice of colors $c\in\{c_{a},c_{b}\}$. To circumvent this issue, we make use of the ancilla qubits in the following way. Let’s consider operators associated to edges of color $c\in\{c_{a},c_{b}\}$. Remember from Lemma 13 that there are exactly $A=r-1$ ancilla qubits. Let’s choose a basis of rainbow subgraphs $R_{1},\ldots,R_{r}$. For each rainbow subgraph $R_{k}$ with $k\leq r-1$, we apply the mapping described above on all the $c$-colored edges but one. On the last edge, we multiply the output operator by an $X$ operator on the $k$th ancilla qubit. For the last edge of the last rainbow $R_{r}$, we transform the input generator by multiplying it by an all-vertex $X$ operator. The output can be obtained by applying the mapping on each $c$-colored edge for which it is already described. Note that all other edges of $\mathcal{G}$ can be written as a product of the edges already described, and the previous discussion therefore completely describes $\mathcal{M}$. An illustration of this is shown in Fig. 18b.
@@
Figure 18: Example of unfolding map on an overlap group. The black $X$s and $Z$s correspond to operators in $O_{A}$, while circled red/blue $X$s and $Z$s correspond to the output operators in $O_{B}$. Following the convention of Ref. [35], we use parentheses when the input is multiplied by elements of the centralizer, which correspond here to an all $X$/$Z$ operator. (a) $Z$ operators on the vertices of every clique are mapped to single-qubit $Z$ operators acting on the clique, except for one operator (the bottom-right in this example), which is only mapped after multiplication by a $Z$ operators acting on all vertices. (b) $X$ operators on edges are mapped to the incident cliques of the opposite color. We look here at red edges only, but the same reasoning applies to blue edges. Dots on top of each graph represent ancilla qubits. The mapping on the first two rows are derived by choosing a basis of $r=3$ rainbow subgraphs (one rainbow subgraph per column) and making use of an ancilla qubit for the last edge of the first $r-1=2$ rainbow subgraphs. For the last edge of the last rainbow subgraph, we need to multiply the input operator by an all-vertex $X$ operator. The ancilla operators of the output are determined by performing this product and using all previously-constructed elements of the mappings. The third and fourth columns can be determined by linearity using elements of the first two columns
@@ B.5 Validity of the mapping
Indeed, each vertex is at the intersection of exactly two $1$-maximal subgraphs, so taking their product gives an identity operator on each vertex. We can show that this is the only relation on $Z$ operators by contradiction. Let’s assume that there exists another relation between the $S_{i}^{Z}$. Let’s consider the subgraph made of all the $1$-maximal subgraphs involved in this relation. Each vertex of this subgraph must be connected to exactly two maximal subgraphs of this set (it must be even due to the relation, non-zero since we are including all the vertices in all the $1$-maximal subgraphs of this subgraph, but cannot be more than two). Therefore, this subgraph must be a $2$-maximal subgraph, which is only possible if it is the full $2$-maximal subgraph. Therefore, the only relation between $Z$ operators is the global relation involving all the $1$-maximal subgraphs. Hence, there are $m-1$ independent $Z$ operators in $O_{A}$
Let’s now take a look at $X$ operators in $O_{A}$. Since those are supported on the two vertices of each edge of the graph, simple cycles gives the relations between them. The number of simple cycles, or circuit rank, of any connected graph is given by $e-v+1$. Therefore, the graph contains $e-(e-v+1)=v-1$ independent $X$ operators.
@@
We now turn to $O_{B}$. It contains single-qubit $Z$ operators for each $1$-maximal subgraph, which are all independent as single-qubit operators. There are therefore exactly $m$ independent $Z$ operators in $O_{B}$. On the other hand, $X$ operators are defined for each edge on the pair of $1$-maximal subgraphs incident to its two vertices. It is easy to check that each cycle of the graph defines a relation on those operators. However, those are not the only relations: every rainbow cycle contains exactly two relations: one for each color of the rainbow cycle, since the product of the operators associated to all edges of the same color in a rainbow cycle is the identity. Therefore, there is a total of $(e-v+1)+r$ relations, and $e-(e-v+1)-r=v-r-1$ independent $X$ operators associated to edges. Adding this up with the number of ancilla qubits (as we have one independent $X$ operator per ancilla), we get:
@@
where we used Lemma 13 to remove $A$ and $r$ from the equation. ∎
@@
In $O_{B}$, every clique supports a single $Z$ operators, so the center does not contain any $X$ operator supported on cliques. However, $X$ operators on the ancilla qubits belong to the center. Therefore, there are exactly $A=r-1$ independent $X$ operators in $Z(O_{B})$. Moreover, $Z$ operators supported on all cliques of a single color also commute with all $X$ operators. Since there are two colors in the flag graphs that we are considering, there are two independent $Z$ operators in $Z(O_{B})$, giving a total number of generators $G(Z(O_{B}))=r+1$. ∎
@@
Let’s choose an $X$ generator $g_{X}\in O_{A}$ and a $Z$ generator $g_{z}\in O_{A}$. Since $g_{X}$ is supported on two vertices, the intersection between $g_{X}$ and $g_{Z}$ must have weight 00, $1$ or $2$.
@@ B.6 Resulting code $C_{B}$
An illustration of such a mapping of stabilizers is shown in Fig. 19
@@ Appendix C Constructing Maximal and Rainbow Subgraphs of Simplex Graphs
[…]
From the simplicial complex, we construct a simplex graph as outlined in II.2 in which the vertices are the $D$-dimensional cells of the complex. Vertices are connected by an edge of colour $c$ if they differ only by a 00-cell of colour $c$.

### v3

@@ Appendix B Unfolding of rainbow codes
An important characteristic of $D$-dimensional colour codes is the possibility to map them into $D$ copies of the toric code through a local Clifford unitary [kubica2015unfolding]. This mapping, often called "unfolding" due its interpretation in two dimensions, has many useful consequences, such as the possibility to construct colour code decoders based on toric code ones [kubica2023efficient] or the understanding of topological excitations and defects in the colour codes [kesselring2018boundaries]
We prove here the existence a local Clifford unitary map between any generic rainbow code constructed from a length-2 chain complex, and the tensor product of two codes whose qubits are supported on the 1-maximal subgraphs of the corresponding flag graph. We call this mapping "unfolding" by analogy with the colour code case, and prove its existence by generalizing the techniques developed in Ref. [kubica2023efficient]. When applying the same toolbox to rainbow codes in higher dimensions or beyond the generic case (e.g. mixed rainbow codes), those techniques seem to fail, which suggests that no such mapping exist in general. We leave this problem open for future work.
@@ B.1 General unfolding process
To prove the existence of an unfolding map, we follow the procedure outlined in Ref. [kubica2015unfolding], that we summarize now. The goal is to construct a local Clifford unitary between a stabilizer code $\mathcal{C}_{A}$ supported on qubits $\mathcal{Q}_{A}$ to a stabilizer code $\mathcal{C}_{B}$ supported on qubits $\mathcal{Q}_{B}$. For this, the idea is to partition the two qubit sets into small parts, $\{Q_{A}^{(i)}\}$ and $\{Q_{B}^{(i)}\}$, and construct a unitary $U_{i}$ for each element of the partition, such that the overall unitary operator can be written as a tensor product of all the $U_{i}$, making it local by construction. To see the effect of such unitary on the stabilizer group of $C_{A}$, we need to consider the overlap of all the stabilizers of $C_{A}$ with each qubit subset $Q_{A}^{(i)}$. If we know how all those stabilizer overlaps are mapped by all the $U_{i}$, we can determine the effect of the overall unitary on the code. The overlap of all the stabilizers of a code $\mathcal{C}$ with any given subset of qubits $Q$ forms a group, that we call the overlap group and write
@@
The following theorem (shown in [kubica2015unfolding]) characterizes an unfolding process:
@@
    $g,g^{\prime}]=[h,h^{\prime}]$ for each $g,g^{\prime}\in\mathcal{G}(O_{A}^{(i)})$, $h=\mathcal{M}_{i}(g)$, $h^{\prime}=\mathcal{M}_{i}(g^{\prime})$.
@@
Note that the first and second conditions prove that the there exists an isomorphism between the two overlap groups, while the last condition ensures that there is unitary implementing the explicit group isomorphism given by $\mathcal{M}_{i}$. The proof of Theorem˜1 heavily relies on the framework constructed in Ref. [yoshida2010framework].
@@ B.2 Outline of the proof
    Proving that this mapping is valid by proving that the three conditions of Theorem˜1 are satisfied.
@@ B.3 Number of ancilla qubits
For the base case, let’s consider a flag graph of clique cardinality $(m_{2},0,0,\ldots)$. The only type of flag graph with no cliques of size higher than $2$ are $2$-rainbow subgraphs. Since $2$-rainbow subgraphs are cycle graphs, the number of vertices is equal to the number of edges, or equivalently to the number of $2$-cliques $m_{2}$. Therefore $v=m_{2}$ and Eq.˜41 is satisfied.
Let’s assume that Eq.˜41 is true for any flag graph of clique cardinality $(m_{2i})_{i\geq 1}$. Let $k\geq 2$. Let’s show that Eq.˜41 is also true for a flag graph $\mathcal{F}$ of clique cardinality $(m_{2},\ldots,m_{2k-2},m_{2k}+1,m_{2k+2},\ldots)$, that is, with one more clique of size $2k$. $\mathcal{F}$ has at least one $2k$-clique of size $2k$. Let’s select one, partition its vertices into $k$ pairs, and delete every edge that does not correspond to a pair of the partition. This new graph, that we call $\widetilde{\mathcal{F}}$, is a valid flag graph, with the same number of vertices as $\mathcal{F}$, but one less $2k$-clique and $k$ more $2$-cliques (corresponding to the edges that have not been deleted). Therefore, $\widetilde{\mathcal{F}}$ has clique cardinality $(m_{2}+k,\ldots,m_{2k-2},m_{2k},m_{2k+2},\ldots)$. By the induction hypothesis, we know that the number of vertices of $\widetilde{\mathcal{F}}$ is given by Eq.˜41:
@@
Since the number of vertices is the same for $\widetilde{\mathcal{F}}$ and $\mathcal{F}$, we can deduce that Eq.˜43 gives the correct identity and Eq.˜41 is verified for $\mathcal{F}$. ∎
@@
Figure 17: Illustration of the proof of Lemma˜14.(a) Flag graph with one 66-clique, one 44-clique, and a rainbow rank $r=4$. (b) Corresponding clique graph, with one node per clique and an edge per connection between two cliques. Rainbow subgraphs of the flag graph correspond to cycles of the clique graph.
@@
To prove this identity, let’s define the clique graph $\mathcal{C}(\mathcal{G})$ associated to $\mathcal{G}$ as a graph where each clique of $\mathcal{G}$ correspond to a vertex in $\mathcal{C}(\mathcal{G})$, and each edge of $\mathcal{G}$ connecting the vertices of two cliques becomes an edge connecting those two cliques in $\mathcal{C}(\mathcal{G})$. An example of clique graph is shown in Fig.˜17. Note that the clique graph is in reality a multigraph, as multiple edges can connect a given pair of nodes. The clique graph has the following useful properties:
@@
Indeed, each $2k$-clique corresponds to a vertex of degree $2k$, and we divide the sum by two to avoid over-counting edges. Inserting this in Eq.˜48, we get
@@ B.4 Mapping construction
The mapping $\mathcal{M}$ is constructed by describing how generators of $O_{A}$ are mapped into operators supported on the cliques of $\mathcal{G}$. For a $Z$ generator supported on the vertices of a clique $\mathcal{C}$, we would like to map it into a single $Z$ operator living on $\mathcal{C}$. However, we note that the product of all the $Z$ generators of $O_{A}$ gives the identity, while the corresponding single-$Z$ operators does not cancel. Therefore, such mapping would map the identity operator into a non-identity operator, and thus cannot be an isomorphism. To solve this issue, we apply the mapping described above to all but one clique. For this clique, we transform the input generator by multiplying it to an all $X$ operator (which is an element of the center of $O_{A}$), and map this new generator to the single $Z$ operator on the clique. You can find an illustration of this in Fig.˜18a.
@@
For $X$ operators, the idea is to map operators supported on the vertices of an edge into an operator acting on the two cliques incident to that edge. However, we have the converse problem as for $Z$ operators: every rainbow subgraph supports two non-trivial $X$ generators that would map to the identity, namely the product of all the $c$-colored edges for each choice of colors $c\in\{c_{a},c_{b}\}$. To circumvent this issue, we make use of the ancilla qubits in the following way. Let’s consider operators associated to edges of color $c\in\{c_{a},c_{b}\}$. Remember from Lemma˜14 that there are exactly A=r1A=r-1 ancilla qubits. Let’s choose a basis of rainbow subgraphs $R_{1},\ldots,R_{r}$. For each rainbow subgraph $R_{k}$ with $k\leq r-1$, we apply the mapping described above on all the $c$-colored edges but one. On the last edge, we multiply the output operator by an $X$ operator on the $k$th ancilla qubit. For the last edge of the last rainbow $R_{r}$, we transform the input generator by multiplying it by an all-vertex $X$ operator. The output can be obtained by applying the mapping on each $c$-colored edge for which it is already described. Note that all other edges of $\mathcal{G}$ can be written as a product of the edges already described, and the previous discussion therefore completely describes $\mathcal{M}$. An illustration of this is shown in Fig.˜18b.
@@
Figure 18: Example of unfolding map on an overlap group. The black $X$s and $Z$s correspond to operators in $O_{A}$, while circled red/blue $X$s and $Z$s correspond to the output operators in $O_{B}$. Following the convention of Ref. [kubica2015unfolding], we use parentheses when the input is multiplied by elements of the centralizer, which correspond here to an all XX/$Z$ operator. (a) $Z$ operators on the vertices of every clique are mapped to single-qubit $Z$ operators acting on the clique, except for one operator (the bottom-right in this example), which is only mapped after multiplication by a $Z$ operators acting on all vertices. (b) $X$ operators on edges are mapped to the incident cliques of the opposite color. We look here at red edges only, but the same reasoning applies to blue edges. Dots on top of each graph represent ancilla qubits. The mapping on the first two rows are derived by choosing a basis of $r=3$ rainbow subgraphs (one rainbow subgraph per column) and making use of an ancilla qubit for the last edge of the first r1=2r-1=2 rainbow subgraphs. For the last edge of the last rainbow subgraph, we need to multiply the input operator by an all-vertex $X$ operator. The ancilla operators of the output are determined by performing this product and using all previously-constructed elements of the mappings. The third and fourth columns can be determined by linearity using elements of the first two columns
@@ B.5 Validity of the mapping
Indeed, each vertex is at the intersection of exactly two $1$-maximal subgraphs, so taking their product gives an identity operator on each vertex. We can show that this is the only relation on $Z$ operators by contradiction. Let’s assume that there exists another relation between the $S_{i}^{Z}$. Let’s consider the subgraph made of all the $1$-maximal subgraphs involved in this relation. Each vertex of this subgraph must be connected to exactly two maximal subgraphs of this set (it must be even due to the relation, non-zero since we are including all the vertices in all the $1$-maximal subgraphs of this subgraph, but cannot be more than two). Therefore, this subgraph must be a $2$-maximal subgraph, which is only possible if it is the full $2$-maximal subgraph. Therefore, the only relation between $Z$ operators is the global relation involving all the $1$-maximal subgraphs. Hence, there are m1m-1 independent $Z$ operators in $O_{A}$
Let’s now take a look at $X$ operators in $O_{A}$. Since those are supported on the two vertices of each edge of the graph, simple cycles gives the relations between them. The number of simple cycles, or circuit rank, of any connected graph is given by ev+1e-v+1. Therefore, the graph contains e(ev+1)=v1e-(e-v+1)=v-1 independent $X$ operators.
@@
We now turn to $O_{B}$. It contains single-qubit $Z$ operators for each $1$-maximal subgraph, which are all independent as single-qubit operators. There are therefore exactly $m$ independent $Z$ operators in $O_{B}$. On the other hand, $X$ operators are defined for each edge on the pair of $1$-maximal subgraphs incident to its two vertices. It is easy to check that each cycle of the graph defines a relation on those operators. However, those are not the only relations: every rainbow cycle contains exactly two relations: one for each color of the rainbow cycle, since the product of the operators associated to all edges of the same color in a rainbow cycle is the identity. Therefore, there is a total of (ev+1)+r(e-v+1)+r relations, and e(ev+1)r=vr1e-(e-v+1)-r=v-r-1 independent $X$ operators associated to edges. Adding this up with the number of ancilla qubits (as we have one independent $X$ operator per ancilla), we get:
@@
where we used Lemma˜14 to remove $A$ and $r$ from the equation. ∎
@@
In $O_{B}$, every clique supports a single $Z$ operators, so the center does not contain any $X$ operator supported on cliques. However, $X$ operators on the ancilla qubits belong to the center. Therefore, there are exactly A=r1A=r-1 independent $X$ operators in $Z(O_{B})$. Moreover, $Z$ operators supported on all cliques of a single color also commute with all $X$ operators. Since there are two colors in the flag graphs that we are considering, there are two independent $Z$ operators in $Z(O_{B})$, giving a total number of generators $G(Z(O_{B}))=r+1$. ∎
@@
Let’s choose an $X$ generator $g_{X}\in O_{A}$ and a $Z$ generator $g_{z}\in O_{A}$. Since $g_{X}$ is supported on two vertices, the intersection between $g_{X}$ and $g_{Z}$ must have weight 0, 11 or $2$.
@@ B.6 Resulting code $C_{B}$
An illustration of such a mapping of stabilizers is shown in Fig.˜19
@@ Appendix C Constructing Maximal and Rainbow Subgraphs of Simplex Graphs
In this Appendix, we set out algorithms for constructing maximal and rainbow subgraphs that are used to define check matrices for rainbow codes. We assume that we are given a simplicial complex of dimension $D$ and that each 0-cell has been allocated a colour from the range $[0..D]$.
From the simplicial complex, we construct a simplex graph as outlined in II.2 in which the vertices are the $D$-dimensional cells of the complex. Vertices are connected by an edge of colour $c$ if they differ only by a 0-cell of colour $c$.